Presentation is loading. Please wait.

Presentation is loading. Please wait.

Status of the LHCb experiment Andrei Golutvin ( Imperial & ITEP & CERN ) on behalf of the LHCb collaboration Outline: Introduction Detector Subsystems.

Similar presentations

Presentation on theme: "Status of the LHCb experiment Andrei Golutvin ( Imperial & ITEP & CERN ) on behalf of the LHCb collaboration Outline: Introduction Detector Subsystems."— Presentation transcript:

1 Status of the LHCb experiment Andrei Golutvin ( Imperial & ITEP & CERN ) on behalf of the LHCb collaboration Outline: Introduction Detector Subsystems Experimental Area Commissioning Cost and funding issues Collaboration matters RRB April 2009 1

2 2 List of the 2008-2009 shut-down activities is nearly completed Full Experimental System Test (FEST) started in December 2008 Exercise Online and Offline systems, Trigger, Monitoring, Data Quality checking and prompt (online) reconstruction Strategy for trigger and subsequent physics analysis of 2009-2010 Run is being prepared LHCb is fully operational for Physics Run RRB April 2009

3 3 Beam Pipe (CERN) 23 m long beam pipe consisting of 3 Be and 1 stainless steel sections Current identified leak fixed by varnishing. The order for a replacement of the third sub-section (UX85/3) was placed, the delivery is expected in 2010 Development of beam pipe supports with reduced material budget Magnet (CF) Warm dipole magnet with 4 Tm bending power along the beam axis Since September 2008 the magnet has not been switched on Perform magnet test in June 2009 and verify improved position stability of the Trigger Tracker and RICH1

4 VErtex LOcator (CH,DE,GB,NL,US,RU) 2×21 pairs of Si sensors arranged in 2 halves; each pair consists of one sensor with R- and one sensor with -strips 4 RRB April 2009 Improvements since last RRB –Software development –Monitoring –Availability of Spare TELL1 (FPGA) boards –Detector Safety / Interlock System –High rate test is underway –Replacement Modules and Mechanics No outstanding problems SUMMARY :

5 VELO Monitoring Online –Extended Diagnostics All in place –Closing Manager NEW realtime vertex monitoring Offline –Diagnostic at pit 5 fold improvement in speed Enhanced reliability and ease of use 5RRB April 2009

6 VELO TELL1 boards TELL1 (84 for VELO) –Receives the trigger and synchronizes the ADCs to the incoming data. The data is digitized and processed During operation in 2008 a major issue was availability of spare TELL1 boards and their reliability Reliability & Spares –No new TELL1 failures since 8/2008 –Now have ~ 12% spare (10/84) –Additional 3 spares in test beam –Additional 3 used for the spare VELO module production 6 RRB April 2009

7 VELO Interlock system Final hardware safety –Prevents operation in modes that could damage the detector Triggers on temperature, pressure, radiation, cooling Acts on HV, LV, Manual Emergency button –Functional (hot) spare existing –Full system up and tested Detector safety remains of the paramount importance to VELO 7 RRB April 2009

8 VELO Replacement (The VELO modules will require replacement after about 3 years of running at full luminosity, due to radiation damage) Rebuild VELO under way –Delivery Q3 2010 –Modules (started) –Hoods & transport (started) –Bases (started) –Cables(ordered) –Evaporators –Vacuum Flanges(ordered) All hybrids and mechanical components completed ~45% of sensors received Attain planned output rate by June 2009 8RRB April 2009

9 Hybrid 206, first fully bonded complete module with silicon sensors and quartz pitch adaptors. Silicon sensorsQuartz Pitch Adaptors Chips 9 RRB April 2009

10 10 Outer Tracker – OT (CERN,CF,CN,DE,NL,PL) Three stations with each 4 stereo layers of straw tubes 5 mm diameter and 5m length; 55k channels OT Calo Muon The installation phase completed in summer 2008 The focus has been on analyzing cosmic and beam data and preparing the detector to 2009/2010 Run Production and test of FE electronics completed The anti-ageing treatment of detector modules at 40 degrees is carried on. At present the entire C-side and two out of 6 C-Frames on the A-side have been treated. Investigations of the treatment results are ongoing. Final report to collaboration in May 2009 SUMMARY :

11 Outer Tracker Ageing The phenomenon is understood: Rapid gain loss due to Araldite AY103 glue Characteristic croissant shape due to formation of O 3 : Aging downstream is prevented –Confirmed with O 3 -like gas in input: NO 2 prevents ageing Relative gain loss: Module length (cm) Module width ( Gas flow High voltage (V) O 3 concentration at output: 11 RRB April 2009

12 Outer Tracker Ageing: strategy Strategy unchanged: Flushing removes glue-vapors (slowly) –Continous flushing since 1.5 year Warming the detector increases outgassing –Continue warming detector in situ O 2 reduces ageing rate –Decide to add few percent O 2 to counting gas HV training (sometimes) repairs damage Flushing O2O2 HV training Trained half Flushing O2O2 12 RRB April 2009

13 Outer Tracker Ageing: strategy Strategy unchanged: Flushing removes glue-vapors (slowly) –Continous flushing since 1.5 year Warming the detector increases outgassing –Continue warming detector in situ O 2 reduces ageing rate –Decide to add few percent O 2 to counting gas HV training (sometimes) repairs; no damage of the wire surface seen HV training Before After 0% O 2 2.5% O 2 286 mV 319 mV O2O2 Small (~15%) gain loss No effect on drift speed 13 RRB April 2009 After

14 Outer Tracker Ageing: in situ test Heating in situ Heating at 30-40 0 C with elec. blankets Heated all 6 C-frames on C-side Heated 2 C-frames on A-side Monitoring in situ Installed scanning frame in situ: T2-Q13-XU T1-Q13-XU 14 RRB April 2009

15 ST: Trigger Tracker & Inner Tracker (CERN,CH,DE,ES,UA) TT covers area of 1.4 1.2 m 2 ; 4 stereo layers with ladders consisting of 3 or 4 chained Si- sensors with strip pitch 183 micron; 143k channels IT: 3 stations with 4 boxes each arranged around beam pipe; each box has 4 stereo layers x-u-v-x, modules with one or two chained Si-sensors; strip pitch 198 micron; 130k channels RRB April 2009 15 Commissioning of both TT and IT detectors is making good progress Broken bonds for TT: majority of the problems occurred a few weeks after module installation and no further problems developed since last RRB - Slow creeping of the pitch adapters to hybrids - Preparing for the production of spares for the case this spreads to other hybrids Further progress in the development of detector control and data quality monitoring SUMMARY :

16 IT Status Service boxes with faults on backplane replaced Replacement of weak optical links (low power optical transmitters) ongoing. To be completed in April After fixes: ~ 99.8 % detector functional [0.5 % with higher noise] Working Faulty fixed [not tested] Faulty Noisy fixed [not tested] Noisy 16 RRB April 2009

17 TT Status 71 low power optical transmitters successfully replaced [10 more to be done next week ] Majority of remaining problems related to broken bonds Will be fixed before start of this years run One module shows HV problem [breakdown @ 160 V ] > 99 % of the detector working for this years run 17 RRB April 2009

18 TT Broken Bonds Innermost row of bonds between hybrid + Pitch Adaptor broken on 8 hybrids Problem not spread to new modules since October Enough spares in hand to replace broken modules In affected modules # broken bonds increases One module, Kapton + pitch adaptor fixed with araldite, problem does not develop further Bondwire 1 pitch adapter side One being used to establish a repair procedure One placed in burn-in stand + thermally cycled [effect not reproduced] All relevant information on production being collected Offer for new hybrid production obtained: - Waiting for better understanding before order placed Two modules removed and taken to Zurich: 18 RRB April 2009

19 RICH (CERN,CF,GB,IT) RICH1 and RICH2 with 3 radiators covers momentum range 2-100 GeV; RICH1: 5cm aerogel with n=1.03 & 4m 3 C4F10 with n=1.0014; RICH2: 100m 3 CF4 with n=1.0005; ~500 HPD to readout RRB April 2009 19 Both RICH1 and RICH2 are complete and routinely taking continuous laser data Further investigation performed to characterize the effects due to vacuum degradation of some HPDs. Degraded tubes have been replaced or are going to be replaced with spares Installation of the Magnetic Distortion Monitoring System for RICH1 has been completed RICH reconstruction software, data quality monitoring and calibration procedures are well advanced SUMMARY :

20 20 RICH Test pattern Regular array of light spots projected over HPD plane in situ in RICH2 (similar results achieved in RICH1 using motorized stage with LEDs) Nicely uniform response and very low noise for almost all the HPDs. A few show noisy behaviour, peaking around the central axis of the tube Due to degraded vacuum quality in those tubes photoelectrons ionize residual gas Ions then accelerated back to the photocathode producing further p.e. ion feedback RRB April 2009

21 21 Ion feedback Ion feedback rate determined using fraction of large clusters ( 5 hits) Rate measured regularly for HPDs over the last 18 months Most show a linear increase of ion feedback with time, with low gradient The noisy tubes have a higher gradient The bad tubes eventually start to glow, but only after ion feedback rate > 5% Glow light RRB April 2009

22 22 HPD repair Failure of tubes can be accurately predicted from the ion feedback measurements ~ 11 HPDs / year predicted to require replacement over the lifetime of the experiment (i.e. 2% / year) Removed tubes are successfully repaired by DEP- Photonis. Opened, cleaned, photocathode reapplied: ion feedback gradient is low after repair Age (days) RRB April 2009

23 RICH2 – Status Map (before intervention) 8_1819_1019_1209_115 00 8_1278_153 9_29_110 01 8_121 9_109_59 02 8_1079_71 8_179 03 8_112 8_609_114 04 8_1478_177 05 9_127 06 9_1559_1488_1359_181 07 8_1338_1249_189 08 9_208 09 9_1379_128_122 9_2659_69 10 8_1609_1428_1979_56 9_3 11 8_1108_134 8_154 12 8_156 9_146 13 8_1578_178 9_159_1259_2508_176 14 8_1199_104 8_1268_141 15 8_1309_1239_118_108 C0C1C2C3C4C5C6C7C8A0A1A2A3A4A5A6A7A8 27 glowing 10 not glowing IFB>5% already 9 not glowing IFB>5% in 2009 11 not glowing IFB>5% in 2010 57 HPDs 23 RRB April 2009

24 HPD RICH2 replacement summary (intervention march 2009) 24 RRB April 2009

25 25 RICH Alignment & calibration After all corrections, reconstructed light spot positions match precisely with regular ~ 10 mm grid pattern Residual contribution to resolution ~ 0.6 mm << pixel size (2.5 mm) correction of magnetic distortion and misalignment is under control x (mm) y (mm) RRB April 2009

26 Calorimeters (CERN,CF,ES,FR,RO,RU,UA) PS/SPD: 12k scint. tiles readout by WLS; ECAL: 6k shashlik cells; HCAL: TILE Calo, 1.5k channels 26 RRB April 2009 The calorimeter system was operational for data taking in September 2008 Detector elements were aligned in time within 3 ns In order to improve stability of the phototube response and minimize the noise for ECAL the Cockroft Walton bases of 6000 PMT have been modified during winter shut-down. The stability tests using LED monitoring system have demonstrated much improved performance of ECAL The HCAL calibration system with Cs source is used regularly to monitor gain calibration Monitoring system for PS/SPD has been commissioned Serializers for the trigger path have been modified and exchanged successfully SUMMARY :

27 Electromagnetic calorimeter ECAL 6000 Ecal PMT bases modified: Improve PMT noise Improve radiation resistance Improve PMT stability first hour excluded (stabilization) 99% of cells are stable within 1% 1% Modified detector fully operational from mid-March 2009 27RRB April 2009 before: rms = 2.9 after: rms = 1.2

28 Calorimeters calibration HCAL calibration 6 Cesium source runs done Reproducibility : ~0.5% Intercalibration ~2% ECAL calibration Derived from LED signal width Expected pre-calibation gain level ~10% ( precision on HV) 24 cells Adjusted calibrations will come from 1st data analysis 28RRB April 2009

29 Muon (CERN,CF,IT,RU) Arranged in 5 SuperLayers; M1 consists of 12 double triple GEM chambers and 264 MWPCs; M2-M5 consists of 1104 MWPCs 29 RRB April 2009 M2-M5 stations have been fully commissioned using cosmic rays Performance agrees well with the expectations Intense debugging activity to fix a number of problematic readout channels (~4% of the total). All the problems understood and most of them fixed (currently ~0.4% problematic channels) Good performance of the HV system. All M2-M5 chambers are at the nominal voltage Very detailed schedule of the M1 installation has been devised and is being followed strictly in order to complete in time. The tests of installed M1 chambers are progressing well SUMMARY :

30 Commissioning of M2-M5 (using ~250k cosmic ray tracks and 19 TED shots in August 2008) 30RRB April 2009 The TED events were recorded with HV @2.3 kV, well below the MWPC plateau for safety reasons Useful to: –test the DAQ chain –check the front-end channels, especially in the inner regions, poorly illuminated by cosmic rays Several missing channels can be seen in the channel maps debugging/recovery A snapshot of the muon detector illuminated by TED events

31 Status of M2-M5 debugging After the cancellation of the LHC run we started a systematic campaign of debugging to fix the problems left open in M2-M5 stations 30/10/200830/03/2009Today Tot log. channels Bad log. channels fractionBad log. channels fractionBad log. channels fraction Side A 83523904.7%650.8%650.8% Side C 83522573.1%170.2%90.1% Total 167046473.9%820.5%740.4% Fixed 57389% A few chambers with HV problems have been replaced and retested: all OK 31RRB April 2009

32 Performance with cosmic rays Cosmic ray data were collected at a different working point wrt to nominal: –Higher thresholds –no CF4 (lower gain) –HV at 2.5 kV (instead of 2.65 – 2.70 kV ideal working point) Performance is to be compared with that expected @ 2.45 kV Expectations: –σ(t) =5.3-6.0 ns –Cluster size in 20ns = 1.02-1.1 –total eff. >0.99 (in an infinite time window) 32RRB April 2009 Not enough statistics for R1 Time resolution from CR data (ns) 2.45 kV Comparison with test beam data agree very well with cosmic data !!!

33 Status of M1 installation Very tight schedule. Exceptional effort by the groups to provide manpower First milestone met: all chambers of the 1 st layer installed by April 10 Some delay in the testing/debugging of 1 st layer Next milestones: –complete alignment and testing of 1 st layer by May 15 –complete the full installation by 15 th of July Complete test of M1 chambers: gas, LV, HV, connectivity, noise 33RRB April 2009 Today: 50% chambers installed Out of them 78% chambers tested

34 RRB April 2009 34

35 RRB April 2009 35 Trigger & Online & Computing (CERN,CF,CH,DE,FR,ES,GB,IT,NL) Full scale HLT testing in realistic environment using FEST Change of strategy for the last phase of HLT (HLT2) Replace OR of many exclusive channels with more inclusive approach CPU Farm is being completed to allow the full 1 MHz readout at nominal event size. The final upgrade is foreseen for beginning of 2010. The readout network will be completed to full capacity during summer 2009 Completing DIRAC3 commissioning for both production and user analysis activities. Continue development of data access system Regular FEST weeks are essential to test data distribution, reconstruction and data quality checks

36 Global Commissioning (Status of LHCb detector alignment) 36 RRB April 2009 SUMMARY :

37 VELO alignment with TED data (TED tracks perfect for VELO alignment: cross detector almost parallel to z) Translation in X of VELO modules extracted from twoTED runs - good agreement between two runs (change with respect to survey less than 20 micron) resolution in μ pitch in μ z of module [cm] Resolution estimated from VELO hit residuals agrees well with expectations Further improvement possible 37 RRB April 2009 R stripsΦ strips

38 IT Alignment Studies used data from the LHC synchronization tests - occupancy 20 times that of normal LHCb running - ghost rate ~ 10 % [not ideal for alignment] Alignment down to granularity of ladders performe d Bias 30 m Core ~120 m Bias 3 m Core 96 m Align Survey - pre-alignment in x Box alignment TxTyRz Layer TxRz, Ladder Tx Tail due to ghosts (For 10 GeV tracks expect 72 m) Unbiased residuals 38 RRB April 2009 Typical example layer IT2X1:

39 TT-Velo and TT-IT alignment TED tracks in IT and VELO can also be matched to TT Residuals of TT hits with respect to tracks in the IT before and after internal alignment of IT remaining shift is possibly due to an IT shearing Residuals of TT hits with respect to tracks in the VELO before and after a global alignment of TT using TT ladder survey not using TT ladder survey next step: relative alignment of 3 systems using 'long' tracks before after 39 RRB April 2009

40 FEST and express stream (VELO alignment / monitoring procedure tested as part of FEST) Set values of velo positioning system such that data 'looks' misaligned Use express stream data to extract new alignment constants Final aim: Use calibration and alignment farm to run alignment within hours after data taking Update alignment before GRID event processing if monitoring tells that old alignment was not good enough difference in X coordinate of primary vertex reconstructed in left and right VELO as seen by monitoring before and after alignment with express stream 40 RRB April 2009

41 41 Cost and Funding Overall cost remains unchanged at 75 MCHF The underfunding in 2005 of 2.6 MCHF for the DAQ CPU Farm, has been covered through extra contributions (BMBF Brazil, France, MPI Germany, Spain, UK, US and CERN) Collaboration Matters Two new groups from UK accepted: - University of Manchester led by Prof. David Bailey - University of Warwick led by Prof. Tim Gershon Service contribution of both groups was extensively discussed and well defined

42 RRB April 2009 42 LHCb Collaboration today: In total: 702 members 15 countries 52 institutes

43 RRB April 2009 43 Possible Physics with 2009/2010 data With 0.3 fb -1 LHCb should be able to improve Tevatron results (expected for 9 fb -1 ) on key observables in flavour physics: 90% C.L. exclusion limits at 8 TeV CM BR(B s 0 μ + μ - ) (x10 -9 ) Mixing phase ϕ s BR(B s μμ)

44 RRB April 2009 44 Conclusion LHCb is well prepared for the 2009/2010 Physics Run We need luminosity in order to be competitive with Tevatron

Download ppt "Status of the LHCb experiment Andrei Golutvin ( Imperial & ITEP & CERN ) on behalf of the LHCb collaboration Outline: Introduction Detector Subsystems."

Similar presentations

Ads by Google